# Search for an Anomalous Production of Charged-Current Interactions Without Visible Pions Across Multiple Kinematic Observables in MicroBooNE

### v2

@@ Abstract
This Letter presents an investigation of low-energy electron-neutrino interactions in the Fermilab Booster Neutrino Beam by the MicroBooNE experiment, motivated by the excess of electron-neutrino-like events observed by the MiniBooNE experiment. This is the first measurement to use data from all five years of operation of the MicroBooNE experiment, corresponding to an exposure of protons on target, a 70% increase on past results. Two samples of electron neutrino interactions without visible pions are used, one with visible protons and one without any visible protons. MicroBooNE data is compared to two empirical models that modify the predicted rate of electron-neutrino interactions in different variables in the simulation to match the unfolded MiniBooNE low energy excess. In the first model, this unfolding is performed as a function of electron neutrino energy, while the second model aims to match the observed shower energy and angle distributions of the MiniBooNE excess. This measurement excludes an electron-like interpretation of the MiniBooNE excess based on these models at CLs in all kinematic variables.
@@
Sideband constraint.—From the set of variations of all the systematic uncertainties under consideration, we calculate the matrix describing the covariances between the predicted event rates in the histograms of the two signal channels and three control channels described above. We then use the block matrix method [32] to constrain the prediction in the 1N0 and 100 signal channels using the data in the and control channels. The constraint is applied to LEE and background predictions, where the LEE is treated in the same way as the electron-neutrinos selected. On average, the constraint procedure reduces the uncertainty on the predicted event rates by 40%. The predictions in the 1N0 and 100 channels are mostly increased after the application of this constraint due to positive correlations between the and predictions and the underpredictions shown in Fig. 1. The full covariance matrix and the impact of these constraints on the predicted event rate and systematic uncertainty are illustrated in the supplemental material [31].
@@
The results of the signal strength fits for all three measured variables, and for each individual signal channel as well as their combination, are shown in Fig. 5, where we have used the Feldman-Cousins [36] procedure to calculate frequentist confidence intervals. When fitting to both channels, the LEE hypothesis is excluded at confidence level (C.L.) in all three variables. The two-hypothesis tests and the signal strength fits for measurements using only the 100 signal channel show a weaker preference for over . This demonstrates that the strong exclusion of the LEE hypothesis in the combined fit is mostly driven by the 1N0 channel.
Summary.—This Letter presents an updated investigation of low energy electron-like events from the BNB using an expanded MicroBooNE dataset of POT. The analysis builds on earlier measurements which test the hypothesis that an anomalous excess is due to an energy-dependent increase in the intrinsic rate and complements this with new tests motivated by the observable kinematics (shower angle and energy) reported by MiniBooNE. In all kinematic variables we find that our data is not consistent with an excess of interactions. Both signal models tested in this measurement that interpret the MiniBooNE LEE as events are excluded at 99% CLs. While the MiniBooNE excess remains unexplained, our measurements confirm and strengthen our previous results. By exploring a more comprehensive set of variables to test the LEE and leveraging an increased statistics dataset, we find that MicroBooNE data is inconsistent with an electron-like interpretation of the MiniBooNE LEE.

### v3

@@ Abstract
This Letter presents an investigation of low-energy electron-neutrino interactions in the Fermilab Booster Neutrino Beam by the MicroBooNE experiment, motivated by the excess of electron-neutrino-like events observed by the MiniBooNE experiment. This is the first measurement to use data from all five years of operation of the MicroBooNE experiment, corresponding to an exposure of protons on target, a 70% increase on past results. Two samples of electron neutrino interactions without visible pions are used, one with visible protons and one without any visible protons. The MicroBooNE data show reasonable agreement with the nominal prediction, with -values when the two samples are combined, though the prediction exceeds the data in limited regions of phase space. The data is further compared to two empirical models that modify the predicted rate of electron-neutrino interactions in different variables in the simulation to match the unfolded MiniBooNE low energy excess. In the first model, this unfolding is performed as a function of electron neutrino energy, while the second model aims to match the observed shower energy and angle distributions of the MiniBooNE excess. This measurement excludes an electron-like interpretation of the MiniBooNE excess based on these models at CLs in all kinematic variables.
@@
Sideband constraint.—From the set of variations of all the systematic uncertainties under consideration, we calculate the matrix describing the covariances between the predicted event rates in the histograms of the two signal channels and three control channels described above. We then use the block matrix method [32] to constrain the prediction in the 1N0 and 100 signal channels using the data in the and control channels. This mathematical procedure modifies the prediction in the signal channels based on the bin-to-bin difference between the data and predictions in the control channels and the uncertainties and correlations across the signal and control channels. The constraint is applied to LEE and background predictions, where the LEE is treated in the same way as the electron-neutrinos selected. On average, the constraint procedure reduces the uncertainty on the predicted event rates by 40%. The predictions in the 1N0 and 100 channels are mostly increased after the application of this constraint due to positive correlations between the and predictions and the underpredictions shown in Fig. 1. The full covariance matrix and the impact of these constraints on the predicted event rate and systematic uncertainty are illustrated in the supplemental material [31].
@@
The results of the signal strength fits for all three measured variables, and for each individual signal channel as well as their combination, are shown in Fig. 5, where we have used the Feldman-Cousins [36] procedure to calculate frequentist confidence intervals. When fitting to both channels, the LEE hypothesis is excluded at confidence level (C.L.) in all three variables. The two-hypothesis tests and the signal strength fits for measurements using only the 100 signal channel show a weaker preference for over . This demonstrates that the strong exclusion of the LEE hypothesis in the combined fit is mostly driven by the 1N0 channel. These results are limited by statistical uncertainties; while this measurement includes data from all MicroBooNE run periods, reconstruction and analysis updates could improve the efficiency and reduce the statistical uncertainty in future electron neutrino analyses.
Summary.—This Letter presents an updated investigation of low energy electron-like events from the BNB using an expanded MicroBooNE dataset of POT. The analysis builds on earlier measurements which test the hypothesis that an anomalous excess is due to an energy-dependent increase in the intrinsic rate and complements this with new tests motivated by the observable kinematics (shower angle and energy) reported by MiniBooNE. For 1N0 and 100 events, in all kinematic variables, we find that our data is not consistent with an excess of interactions, with -values showing agreement with the null hypothesis in the 10–77 range, depending on the final state and the kinematic variable considered. Both signal models tested in this measurement that interpret the MiniBooNE LEE as events are excluded at 99% CLs. While the MiniBooNE excess remains unexplained, our measurements confirm and strengthen our previous results. By exploring a more comprehensive set of variables to test the LEE and leveraging an increased statistics dataset, we find that MicroBooNE data is inconsistent with an electron-like interpretation of the MiniBooNE LEE.